Dr. Santiago and two of his colleagues are considering opening a new urgent care clinic in a large metropolitan area. Their intent is to provide easy access for patients by having the clinic open 360 days per year, 16 hours each day from 7 a.m. to 11 p.m. The urgent care center would be staffed by a physician, medical assistant, and patient service coordinator for each of the two 8-hour shifts.

In order to determine the feasibility of the project, Dr. Santiago hired a business development consultant to assist with market projections. The results of the study show that if the physicians spent $450,000 on advertising the first year, they could expect about 75 new patients each day. Dr. Santiago and his associates believe the estimate to be reasonable and are prepared to spend the $450,000 on advertising.

Assume the urgent care center receives $80 for the initial consultation. For patients who require additional services, the center will receive an extra $40 for the visit. Dr. Santiago estimates that 35% of consultations will result in additional services.

The hourly wages of the staff are projected to be $100 for the physician, $18 for the medical assistant, and $14 for the patient service coordinator. Fringe benefit expense will be 40% of the wages paid.

Dr. Santiago has located 6,000 square feet of suitable clinic space for the new venture, which rents for $30 per square foot annually. Associated expenses will be $12,000 for property insurance and $30,000 for utilities (phone, internet, electricity, waste disposal and cleaning). The group will purchase malpractice insurance for $120,000 annually. The initial investment in medical and office equipment will be $75,000. The cost of supplies will be $8 for initial consultation and $15 total for patients who require additional services.

1) Determine the total net income the new urgent care center can expect if all goes according to plan in its first year of operations.

2) Determine how many patients must visit the urgent care center per day for the venture to break even during its first year of operation.

I will thumbs up all answers! Thanks so much!
Answer #1
Patients 27,000
Patients needing additional Services 9,450
Revenue from Initial Consultation $        2,160,000
Revenue from Additional Services $            378,000
Total Revenue $        2,538,000
Less - Variable Cost
Cost of Supplies
Initial Consultation $            216,000
Additional Services $              66,150
Total Variable cost $            282,150
Contribution $        2,255,850
Contribution Margin 89%
Less - Fixed Cost
Salaries
Physician $            576,000
Medical assistant $            103,680
Patient service coordinator $              80,640
Fringe benefit $            304,128
Rent $            180,000
Associated expenses $              12,000
Total Fixed Expenses $        1,256,448
Net Income $            999,402

Break even in no. of patients =

Contribution from intial consultation = 80 - 8 = $72

Contribution from addtional service = 120 - 15 = $105

Weighted contribution = 72* 1 + 105*0.35 = 109

Breakeven = 12,56,448 / 109 = 32

Patient per day = 32/100*135 * 360 days = 43 patients per day